A car starts motion from rest and travels for 5 sec. with a uniform acceleration of +1.5 m/sec2. The driver then applied the brakes, causing a uniform acceleration of -2 m/sec2 if the brakes are applied for 3 sec. How long it takes the car from the beginning of the motion to get to a complete stop?


1
Expert's answer
2021-06-10T16:42:10-0400
"0.5a_1t_1^2=0.5a_2t_2^2\\1.5(5)^2=2t_2^2\\t_2=4.33\ s"